OpenStudy (anonymous):

Log(-x-1)=log(5x)-logx When I solve for x I got -6 but the book answer says no solution. Can anyone show me why? @precal

OpenStudy (precal):

definition of log

OpenStudy (precal):

x can't be negative

jimthompson5910 (jim_thompson5910):

if x = -6, then log(-x-1) = log(5x) - log(x) log(-(-6)-1) = log(5(-6)) - log(-6) log(5) = log(-30) - log(-6) But you can't take the log of a negative number, so x = -6 isn't part of the domain

OpenStudy (anonymous):

But shouldn't it have said -6 is extraneous? I am concerned that I did the solving steps incorrectly.

OpenStudy (precal):

no solution and extraneous can be used the same way........